you run a manufacting business that is involved in manufacturing and selling a single product. The annual fixed expenses to run the business are 15000$ and variable expenses are 7.50$ per unit.The sale price of your pouduct is 15$ per unit.calculate(a) the number of units to be sold to breakeven point,(b)the breakeven point in dollars
Answer #1
FIXED COST 15000
VARIABLE EXPENSES per unit 7.5
SELLING PRICE PER UNIT 15
Let x be the breakeven number of units
15*x = 7.5*x + 15000
7.5*x = 15000
x = 2000
a) The number of units to be sold to breakeven point is 2000.
SALES REVENUE 15*2000
SALES REVENUE $30000
b) The breakeven point in dollars is equal to $30000.